Is (7,4) on the graph of f(x) = 3x+6/2x-3 ? Explain!

Answer by edjones(8007)

f(x) = (3x+6)/(2x-3) I assume that this is the equation
If it is then (3x+6)/(2x-3) will have to equal 4 when x=7
(3x+6)/(2x-3)
=((3*7)+6)/((2*7)-3)
=(21+6)/(14-3)
=27/11
No
.
Ed